amourniyaa05:

what is (5n+6xn)-(8n9)

sophia333:

Reorder your terms.

sophia333:

Then solve variable "n"

amourniyaa05:

ugh so hard

sophia333:

Dm me. I can help you one on one. Okay?

sophia333:

Or i could just help you here-

sophia333:

Do you know how to simplify 6-8n=5n+19?

sophia333:

6 + -8n = 5n + 19 this is it simplified.

sophia333:

So now reorder your terms as i said.

sophia333:

Do you know how to reorder?

sophia333:

Reordering it would be 6 + -8n = 19 + 5n

sophia333:

Move all terms containing n to the left, all other terms to the right.

sophia333:

Add "-5n" to each side of the equation. 6 + -8n + -5n = 19 + 5n + -5n

sophia333:

Combine alike terms: -8n + -5n = -13n

sophia333:

6 + -13n = 19 + 5n + -5n

sophia333:

5n + -5n = 0

sophia333:

6 + -13n = 19 + 0

sophia333:

6 + -13n = 19

sophia333:

Add "-6" to each side of the equation. 6 + -6 + -13n = 19 + -6

sophia333:

Combine alike terms: 6 + -6 = 0 0 + -13n = 19 + -6 -13n = 19 + -6

sophia333:

19 + -6 = 13

sophia333:

-13n = 13

sophia333:

Divide each side by "-13". n = -1

sophia333:

So your answer is?

sophia333:

@amourniyaa05 are you still there?

amourniyaa05:

yes im here

sophia333:

So what would your answer be?

amourniyaa05:

idk atm

sophia333:

What did i say you needed to find?

sophia333:

N, correct?

amourniyaa05:

idk

sophia333:

Please cooperate with me. \(\color{#0cbb34}{\text{Originally Posted by}}\) @sophia333 Then solve variable "n" \(\color{#0cbb34}{\text{End of Quote}}\)

amourniyaa05:

okayy

sophia333:

So what is variable n equivalent to?

sophia333:

I just gave you all the work you need. Just reread.

amourniyaa05:

no

sophia333:

Okay well, i cannot help you unless you cooperate. Therefore, you won't have your answer. QuestionCove has a cheating policy. As I said all you need is to reread what I said variable "n" was.

sophia333:

If you don't believe me here's the rules https://questioncove.com/rules If you really needed help all you would do is scroll up :)

amourniyaa05:

okay im ready ii was tryna work it out.

sophia333:

So, now that you worked it out, what do you think the answer is?

sophia333:

I gave it to you, so it shouldn't be hard.

sophia333:

@amourniyaa05 Its not hard.

sophia333:

Divide each side by "-13". n = -1

amourniyaa05:

to me it is.

sophia333:

So what is n?

sophia333:

Its not hard because i gave you the answer

sophia333:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sophia333 Divide each side by "-13". n = -1 \(\color{#0cbb34}{\text{End of Quote}}\) Cmon- its right there-

amourniyaa05:

um

amourniyaa05:

-1

sophia333:

Correct. That is your answer.

sophia333:

(5n+6xn)-(8n9) is -1.

amourniyaa05:

the answer is -1

sophia333:

Yeah

amourniyaa05:

thanks boo.

sophia333:

No problem